OpenStudy (anonymous):

What is 0.36 expressed as the quotient of two integers in simplest form?

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@lilshane

OpenStudy (anonymous):

how do you write any decimal as a fraction?

OpenStudy (anonymous):

@mick14731 wat do u mean

OpenStudy (ahsome):

@k12user, lets say the number is 36 and 100: \[\Huge{\frac{36}{100}}\] Why 36 and 100? Because then its easy to get any number in decimal form, simply divide by 10, 100, 1000etc Now, lets see if we can simplify these numbers \[\Huge{\frac{36}{100}/2}\] \[\Huge{\frac{18}{500}}\] \[\Huge{\frac{18}{500}/2}\] \[\Huge{\frac{9}{25}}\] Its in its most simplified form. Therefore, the two smallest whole numbers that give 0.36 are 9 and 25 :)

OpenStudy (anonymous):

thx @Ahsome

OpenStudy (ahsome):

No problem, @k12user :D